• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jroche

    Log entry time 00:42:53 on October 29, 2010

    Entry number 337256

    keyword=Helicity flip rate restored to 943 Hz

    M. Friend noticed that the helicity flip rate was set to 120 Hz today at 11 am during beam studies. The main DVCS data taking do not care about the value of the flipping rate but the Compton DAQ does. We decided to revert to the standard setting (in between run 8018 and 8019 for main DVCS DAQ, and 23575 and 23576 for the Compton DAQ). The current setting is: Helicity Delay Set : 8 windows Helicity Pattern Set : Quartet Helicity Clock Set : Free Clock Helicity Settle Read : 60 Helicity Stable Read : 1000 Helicity Delay Read : 8 windows Helicity Pattern : Quartet Helicity Clock Read : Free Clock Helicity Frequency : 943.396